A system absorbs 50J of heat during a transformation. Determine the change in internal energy if 


the system perform, (a) 15J and (b) 65J of work on the surrounding.


Find the ∆E


a) When 175J of work is done on a system that evolves 50J of heat to the surrounding.


b) When a gas absorbs 28kJ of heat and 13kJ of work done on it.

1
Expert's answer
2022-03-27T10:28:34-0400

a. W=-15J

q=50

"\Delta U=q+W=50-15=45 J"

W=-65

"\Delta U=q+W=50-65=-15 J"



a) q=-50

W=175

"\Delta U=q+W=-50+175=125 J"

b) q=28000

W=13000

"\Delta U=q+W=28000+13000=41000 J"